OpenStudy (anonymous):

A triangle is formed by the x-axis, the y-axis and the line y = ax + b. If a = -b, then what is the volume of the cone generated by rotating this triangle around y-axis. a) pi/9 b) (pi/3)a c) pi d)3pi

OpenStudy (anonymous):

b is the height the radius is 1

OpenStudy (anonymous):

umm.. i don't get it :s

OpenStudy (anonymous):

Y = ax + b becomes y = -bx + b y/b = -x + 1 from this we can see that intercept on y axis = 1 and this is the height of the cone. when y = 0, x = 1 and that is the radius of the base of the cone

OpenStudy (anonymous):

sorry secret there is an error in the above the radius of the cone is 1 but i'll take another look at the height

OpenStudy (anonymous):

the height is b as him said - the intercept on the x axis is -b1/2 so the volume is (1/3) pi r^2 h = (1/3) pi 1^2 * -b =(1/3)* pi * a = (pi/3) * a which is b) on the list

OpenStudy (anonymous):

i'm going crazy here correction of the first line the intercept on the y axis is -b!!

OpenStudy (anonymous):

i've got it right at last

OpenStudy (anonymous):

oh, coool. NOW i really understand it. thanks guys!

OpenStudy (anonymous):

yw